OpenStudy (anonymous):

A wall can be built by 6 men working 8 hours per day in 5 days. How many days will it take 4 men to build the wall if they work only 5 hours per day??

OpenStudy (anonymous):

calculate how much work one man can do per hour.

OpenStudy (anonymous):

what did you try?

OpenStudy (anonymous):

i got 1 man = 15/4 hours = 1 day

OpenStudy (anonymous):

yes, but what did you do to get that?

OpenStudy (anonymous):

6 men = 8/5 hours = 1 day then 6/8/5=15/4 1 man = 15/4 hours = 1 days

OpenStudy (anonymous):

i meant day

OpenStudy (anonymous):

okay. you did it all wrong.

OpenStudy (anonymous):

the total work was done by 6 guys in 5 days working 8 hours a day. that means, the total work was done by 6 guys in 40 hours, right?

OpenStudy (anonymous):

so, in one hour, 6 guys can do 1/40th of the work. that means 1 guy can do 1/6th of 1/40th of the total work

OpenStudy (anonymous):

or 1 guy can do 1/240th of the total work in one hour.